Find the value of c for which P(x < c) = 0.7.


1
Expert's answer
2021-05-26T16:25:06-0400

Uniform distribution "U(a, b)"


"x\in [a, b]"

"f(x; a, b) = \begin{cases}
 \dfrac{1}{b-a} &a \leq x \leq b\\
 0 &\text{otherwise } \end{cases}"

For "a \leq c\leq b"


"P(x<c)=\dfrac{c-a}{b-a}"

Given


"P(x<c)=0.7"

"\dfrac{c-a}{b-a}=0.7"

"c=a+0.7(b-a)"

"c=0.3a+0.7b"